OpenStudy (ashlyhall8):

Factor completely, then place the factors in the proper location on the grid. 9 - 6a - 24a2

OpenStudy (instagrammodel):

Set the factor '(1 + -2a)' equal to zero and attempt to solve: Simplifying 1 + -2a = 0 Solving 1 + -2a = 0 Move all terms containing a to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -2a = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -2a = 0 + -1 -2a = 0 + -1 Combine like terms: 0 + -1 = -1 -2a = -1 Divide each side by '-2'. a = 0.5

OpenStudy (instagrammodel):

I think..

OpenStudy (denonakavro):

The person on top is wrong

OpenStudy (instagrammodel):

Oh. Okay.

OpenStudy (denonakavro):

9 - 6a - 24a2 First rearrange terms -24a2 -6a +9 Now factor -3(8x^2+2x-3) By further simplifying the (8x^2+2x-3) we get (-3)(2x-1)(4x+3) as the answer Use this site as reference http://www.purplemath.com/modules/factquad2.htm